OpenStudy (anonymous):

which of the following equals 1-(sin^2feta/1-cosfeta)? a.cos feta b.-cos feta c.1-sin feta d.1+sin feta

OpenStudy (anonymous):

(1-cos theta - sin^2 theta) / (1 - cos theta) = (cos^2 theta - cos theta) / (1 - cos theta) = cos theta